OpenStudy (bloomlocke367):

you place 3.52 g of sodium hydroxide to dissolved in a calorimeter cup. the cup contains 1000 mL of water at 20 degrees celcius (in a calorimeter cup). as the sodium hydroxide dissolves, the temperature of the surrounding water increases to 23.7 degrees celsius. Complete the table below and determine the total energy released.

OpenStudy (bloomlocke367):

@jdoe0001 @whpalmer4

OpenStudy (bloomlocke367):

For the chart I need to know the mass of water, heat capacity of water, change in temperature, and the total energy released.

OpenStudy (bloomlocke367):

the change in temperature is 3.7 degress celsius

OpenStudy (anonymous):

0.352kg * 4.184 * 3.7 = energy needed

OpenStudy (anonymous):

in joules

OpenStudy (bloomlocke367):

is the energy needed the same as the energy released? where'd those numbers come from?

OpenStudy (bloomlocke367):

and for the energy needed, I got 5.4492416

OpenStudy (anonymous):

3.52g=0.00352 Kg 23.7-20.0=3.7 degrees C 4.184 = specific heat of water

OpenStudy (bloomlocke367):

okay. Well my dad said I gotta stop for the night. xc I guess I'll finish tomorrow,,, thanks for the help

OpenStudy (bloomlocke367):

@amistre64 @JFraser

OpenStudy (bloomlocke367):

so is what I have correct so far? I have 1000g for the mass of water, 4.18 J/g°C for the heat capacity of water, the change in temp is 3.7°C, and the total energy released is 5.45 J?

OpenStudy (bloomlocke367):

Next I need to determine the heat of solution of the sodium hydroxide in J/g... how do I do that?

OpenStudy (bloomlocke367):

@JFraser can you help?

OpenStudy (jfraser):

the \(reason\) for the energy release is the dissolving of the sodium hydroxide, which released 5.45J, as you've found. If you started with 3.52g of sodium hydroxide, and it released 5.45J of energy, the J/g is .... \[\frac{5.45J}{3.52g}\]

OpenStudy (bloomlocke367):

ohhhhhhhhh okay. so it's 1.5482954.... rounded to 1.55J/g with the right amount of Sig, Fig. right?

OpenStudy (jfraser):

right

OpenStudy (bloomlocke367):

okay.. that makes sense. can you help with another one? thanks so much for your help!

OpenStudy (lexilove13):

@BloomLocke367 Hey! Do you have it done?